OpenStudy (anonymous):

a certain number of men completed a work in 60 days. if there were 8 men more, the work could be finished in 10 days less. how many man were orignally der ??

OpenStudy (anonymous):

x * 60 = Y (x+8) * 50 = Y 60X = 50x+400 10x = 400 x = 40 So, There were 40 Men worked originally.

OpenStudy (anonymous):

k thx